# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

### v1

@@ 2.2.2. Hierarchical Pooling-based Graph Neural Networks
  • Similarity-based. Similarity-based approaches (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b) cluster nodes based on the similarity of nodes features or structures. These methods rely on predefined similarity metrics or graph partitioning algorithms to generate the cluster assignment matrix. Graclus (Mesquita et al., 2020; Dhillon et al., 2007) and CC-GNN (Li et al., 2022b) assign nodes to different clusters based on feature similarity and graph structure.
@@
2.2.3. Subgraph-based Graph Neural Networks
@@
2.2.4. Graph Learning-based Graph Neural Networks
Due to uncertainty and complexity in data collection, graph structures may be redundant, biased, or noisy. Consequently, GNNs cannot learn reliable and accurate graph representations, leading to incorrect predictions. Therefore, as shown in Figure 1 (d), current researchers (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose reconstructing the graph structure and node features to improve the quality of the learned graph representations. Given a labeled graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, the graph learning-based approaches can be formulated as bi-level optimization problem.
@@ 2.2.4. Graph Learning-based Graph Neural Networks
(11) $s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}),\forall G_{i}\in\mathcal{G}$
@@
  • Preprocessing-based. Preprocessing-based approaches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly. The basic idea is to first reveal the common graph patterns and modify the graph structure and node features to recover these patterns. For example, GNN-Jaccard (Wu et al., 2019) and GNAT (Li et al., 2022a) observe that similar nodes tend to be connected by edges. Therefore, they remove edges between dissimilar nodes and add edges between similar ones. Also, GNN-SVD (Entezari et al., 2020) observes that noisy edges and node features tend to increase the rank of the adjacency matrix. Therefore, GNN-SVD reconstructs graphs by decreasing the rank of the adjacency matrix.
  • Jointly Training-based. In contrast to static preprocessing-based methods, jointly training-based approaches aim to optimize the GNN while iteratively reconstructing the graph structure and node features alongside the GNN model parameters. This iterative bi-level optimization enables the GNN to dynamically adapt the graph structure and features based on task-specific objectives. For instance, several approaches (Jin et al., 2021, 2020b; Li et al., 2024b; Franceschi et al., 2019; Sun et al., 2022; Luo et al., 2021), such as ADGNN (Li et al., 2024b), ProGNN (Jin et al., 2020b), and SimPGCN (Jin et al., 2021), propose reconstructing the edges between each pair of nodes by minimizing the GNN loss on the reconstructed structure while also reducing the rank of the adjacency matrix. Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
2.2.5. Self Supervised Learning-based Graph Neural Networks
@@
Depending on the technique of generating supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main types.
  • Pretext Task-based. Pretext task-based approaches (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020a; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels. Current approaches design diverse node-level and graph-level tasks to capture intrinsic graph patterns (Wang et al., 2022b), such as predicting node attribute, node degree, and the node number. For example, HMGNN (Zang et al., 2023) pretrains GNNs by predicting the links between nodes and the number of nodes in each graph. MGSSL (Zhang et al., 2021) first masks the edges among motifs and then predicts the missing edges. MoAMa (Inae et al., 2023) first masks node features and then uses GNNs to reconstruct the node features based on the masked graphs. GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020a) predict both node attributes and edges.
  • Graph Contrastive Learning-based. Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples. The basic idea is to maximize the similarity between different views or augmentations of the same node or graph (positive pairs) while minimizing the similarity between different nodes or graphs (negative pairs). In such a way, GCL-based approaches can distinguish the similarity between nodes and graphs. In general, the SSL loss $\mathcal{L}_{ssl}(\cdot)$ in Equation (12) can be formulated as follows.
@@
    (15) $s.t.$ $\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
    where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss. Additionally, $\mathcal{L}_{positive}(\cdot)$ is the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views. Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs with labels different from $G_{i}$. One typical $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (14) can be defined based on InfoNCE loss (Zhu et al., 2021; Yeh et al., 2022) as follows:
@@
    The techniques in Equation (15) are employed to generate positive views for each node. Specifically, similarity-based methods (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes. Also, diffusion-based approaches (Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) modify the graph’s structure (e.g., by deleting or adding edges) based on global topological information, such as personalized PageRank (PPR) (Haveliwala, 1999) and motif-preserving (Sun et al., 2024). Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu et al., 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.
3. A Generalizable and Expressive GNN
We propose a simple yet effective contrastive learning-based GNN with high generalization and expressivity capabilities on graph-level tasks. As shown in Figure 1, our proposed GNN model consists of three components, i.e., adaptive graph augmentation, $k$-Path rooted subgraph encoder, and consistency-aware model training.
Figure 2. Framework overview.
3.1. Adaptive Graph Augmentation
To enhance the generalization ability of the GNN model, we propose using contrastive learning to optimize GNNs for graph-level tasks. As described in Section 2, the key step is to generate positive graph pairs for each graph. Here, we propose a unified contrastive learning approach for graphs in diverse domains, such as geometric graphs and biology. The basic idea is to first measure the importance of each edge concerning the whole graph and then drop the unimportant edges based on weighted sampling to generate positive pairs for $G$.
Specifically, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ with edges $E$, the edge importance of each edge $e_{v,u}\in E_{i}$ is calculated based on node degree of $v$ and $u$ (Li et al., 2024a; Zhu et al., 2020), i.e., $imp_{v,u}=\log\left(\frac{D_{i}(v)+D_{i}(u)}{2}+1\right)$, where edge connected with higher-degree nodes tends to be important because nodes with higher degrees in a graph are bridges and core nodes in a functional motif as follows:
  • Geometric Graphs. High-degree nodes act as crucial bridges, and removing their edges significantly impacts the number of paths and cycles within this graph.
  • Social Networks. Higher-degree nodes are crucial as they often represent key influencers or hubs in the network. These nodes connect different parts of the network, facilitating communication and information flow. Their removal can fragment the network and disrupt connectivity significantly.
  • Chemistry Graphs. High-degree atoms mainly determine the properties of a motif. For example, phosphorus (P) in phosphate groups ($\text{PO}_{4}$) is bonded to four oxygen atoms (O). While oxygen atoms contribute to properties like electronegativity and polarity, phosphorus is often more critical in determining the overall functionality and behavior of the phosphate group.
Then, the dropping probability $p_{v,u}$ of each edge $e_{v,u}$ can be normalized using min-max normalization (Zhu et al., 2020; Patro, 2015; Li et al., 2024a) as follows:
(17) $p_{{v,u}}=\mu\cdot\frac{imp_{max}-imp_{v,u}}{imp_{max}-imp_{min}},$
where $\mu\in[0,1]$ is a hyperparameter, and $imp_{max}=\max_{e_{i,j}\in E_{i}}imp_{i,j}$ and $imp_{min}=\min_{e_{i,j}\in E_{i}}imp_{i,j}$ are the largest and smallest edge score, respectively. Then, for each graph $G_{i}$, we can generate two positive graphs $\hat{G}_{i}$ and $\tilde{G}_{i}$ by randomly dropping unimportant edges based on Equation (17).
3.2. $k$-Path Rooted Subgraph Encoder
A more expressive GNN (Xu et al., 2019; Huang et al., 2022). can better encode substructures like paths, cycles, and other motifs, which are essential for solving downstream tasks. To enhance the expressivity ability, we propose a $k$-Path rooted subgraph encoder following (Huang et al., 2022; You et al., 2021).
3.2.1. $k$-Path Rooted Subgraph and Feature Augmentation
Given a graph $G(V,\mathbf{A},\mathbf{X})$, we define a $k$-tuple set of each node $v_{i}\in V$ as $\mathcal{T}_{v_{i}}^{k}=\{\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})\}$, where there is an edge between each consecutive pair $(v_{i},v_{i+1})$ and $v_{i+j}$ is actually a $j$-hop neighbor of node $v_{i}$. Particularly, for each node $v_{i}\in V$, the number of $k$-tuples starting from $v_{i}$ is $O(\overline{n}_{nei}^{k})$, where $\overline{n}_{nei}$ is the average neighbor number of nodes in $G$. Then, for each tuple in $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$, we can extract a subgraph based on nodes in $\vec{t}_{v_{i}}^{k}$. Formally, given a $k$-tuple nodes $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ and the graph $G$, we denote the rooted subgraph of $\vec{t}_{v_{i}}^{k}$ as $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$, which extracts $L$-hop neighbors of each node in $\vec{t}_{v_{i}}^{k}$, i.e., $V^{s}_{\vec{t}_{v_{i}}^{k}}=\cup_{v_{j}\in\vec{t}_{v_{i}}^{k}}{\cup_{l=0}^{L}\mathcal{N}^{l}(v_{j})}$ and keep the edges among nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ in $G$.
Then, given each rooted subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$ of nodes $\vec{t}_{v_{i}}^{k}$, we add identifiers for each node in $\vec{t}_{v_{i}}^{k}$ based on the consecutive order. Specifically, for each node $v_{i+j}\in\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{i+j})\in\mathbb{R}^{d_{x}}$ is augmented as $\mathbf{x}(v_{i+j})=\mathbf{x}(v_{i+j})\oplus\mathbf{e}_{j,k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{e}_{j,k}\in\{0,1\}^{k}$ is a one hot vector, where the $j$-th element is 1 and others are 0. For nodes $v_{m}\in V^{s}_{\vec{t}_{v_{i}}^{k}}\setminus\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{m})\in\mathbb{R}^{d_{x}}$ of is augmented as $\mathbf{x}(v_{m})=\mathbf{x}(v_{m})\oplus\mathbf{0}_{k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{0}_{k}$ is a vector of length $k$ and all its components are 0. Then, we denote the feature augmented subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}$ as $G^{s,a}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s,a}_{\vec{t}_{v_{i}}^{k}})$.
3.2.2. Graph Representation Learning
Given each augmented rooted subgraph $G^{s,a}_{\vec{t}_{v_{i}}^{k}}$, the node representations of nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ can be learned by a GNN encoder as $\mathbf{H}({V^{s}_{\vec{t}_{v_{i}}^{k}}})$, such as GraphSAGE (Hamilton et al., 2017) or GatedGNN (Li et al., 2016). Then, the representations of each node $v_{i}$ can be learned based on all subgraphs of each $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ as follows:
(18) $\mathbf{h}({v_{i}})=\mathsf{READOUT}_{1}(\{\mathbf{H}(V^{s}_{\vec{t}_{v_{i}}^{k}})[v_{i}]|\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}\}),$
Where $\mathsf{READOUT}_{1}(\cdot)$ is a readout function, such as MaxPooling. Therefore, the final graph representation of $G(V,\mathbf{A},\mathbf{X})$ can be learned as follows:
(19) $\mathbf{h}=\mathsf{READOUT}_{2}(\{\mathbf{h}({v_{i}})|v_{i}\in V\}),$
The graph representation $\mathbf{h}$ can be used to predict graph labels or graph properties, as mentioned in Equation (1).
3.2.3. Theoretical Analysis
Theorem 3.1.
Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+2$-Path, starting from the node $v_{i}$.
Proof.
We firstly give Lemma 3.2 and then prove Theorem 3.1.
Lemma 3.2.
Given a graph denoted as $G(V,\mathbf{A},\mathbf{X})$ and a $k$-tuple $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ starting from node $v_{i}$, our model can count the number of $(k+2)$-Path between $v_{i}$ and any node $v_{j}\notin\vec{t}^{k}_{v_{i}}$ that pass through the path $(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$.
Proof.
Note that since we extract the subgraphs based on the $k$ tuple $\vec{t}_{v_{i}}^{k}$, we can know whether a node belong to this $\vec{t}_{v_{i}}^{k}$. Given a node $v_{j}\in V$ and $v\notin\vec{t}^{k}_{v_{i}}$, we can count the $k+2$-Path between $v_{i}$ and $v_{j}$ by aggregating information from neighbor. Firstly, we can identify whether a node $v_{j}$ is the one-hop neighbor of the last element of $\vec{t}_{v_{i}}^{k}$ (i.e., $v_{i+k-1}$) as follows:
(20) $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}}[1:k-1])\sum_{v_{m}\in N(v_{j})}{\mathbb{I}(v_{m}=v_{i+k-1})}$
where $\vec{t}_{v_{i}}^{k}[1:k-1]=(v_{i},v_{i+1},\cdots,v_{i+k-2})$ is the first $k-1$ elements of $\vec{t}_{v_{i}}^{k}$, the indicator function $\mathbb{I}(condition)=1$ if the $condition$ is true. Specifically, in Equation (20), if $v_{j}$ is the one-hop neighbor $v_{i+k-1}$, the hidden representation $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=1$ otherwise $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=0$.
Then, in Equation (21), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 2-Path from node $v_{i+k-1}$ to the node $v_{j}$.
(21) $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{m})}$
Finally, in Equation (22), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 3-Path from node $v_{i+k-1}$ to the node $v_{j}$.
(22) $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{(\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j}))},,$
If $v_{j}$ is a one-hop neighbor of $v_{i+k-1}$, there exists a recursive 3-Path of the form $(v_{i+k-1},v_{j},v_{m},v_{j})$. To eliminate these recursive 3-Path for $v_{j}$, we use $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$.
Since $\vec{t}^{k}_{v_{i}}$ is a $k-1$ path from node $v_{i}$ to node $v_{i+k-1}$ and $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of 3-Path from $v_{i+k-1}$ to $v_{j}$, $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of $k+2$-Path from $v_{i}$ to $v_{j}$ that pass through the path $\vec{t}^{k}_{v_{i}}$.
Based on Lemma 3.2, given the graph $G(V,\mathbf{A},\mathbf{X})$, we can compute the total number of $k+2$ path in the graph $G$ that starts from node $v_{i}$ as follows: $n^{k+2}_{v_{i}}=\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\sum_{v_{j}\in V}{\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})}.$
Theorem 3.3.
Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+3$-Cycle of the node $v_{i}$.
Proof.
Based on the Theorem 3.1, we can counting the number of $k+2$-Path from each node $v_{i}$ that passes through the path $\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}$, then we can cunt the number of $k+3$-Cycle of node $v_{i}$ based on its neighbor $N(v_{i})$ as $\sum_{v_{j}\in N(v_{i})}\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\mathbf{h}^{(3)}_{\vec{t}^{k}_{v_{i}}}$. ∎
3.3. Consistency-aware Model Training
We propose a contrastive loss and a task-aware consistency loss to optimize GNNs, enhancing the generalizability of learned representations and effectively capturing graph patterns. Given a batch of graphs $\mathcal{G}_{B}=\{(G_{i},y_{i})\}_{i=1}^{B}$ with size $B$, we generate two positive views $\hat{G}_{i}$ and $\tilde{G}_{i}$ for each $G_{i}\in\mathcal{G}_{B}$ by dropping edges from $G_{i}$ based on Equation (17). The contrastive loss $\mathcal{L}_{cl}(f_{\theta},G_{i})$ for each graph $G_{i}$ is defined as follows:
$-\log\frac{\sum_{\{G_{j}|y_{j}=y_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{j}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s({\mathbf{h}}^{\prime}_{i},{\mathbf{h}}^{\prime}_{j})+s(\hat{\mathbf{h}}_{i},\mathbf{\tilde{h}}_{i})}}{\sum_{G_{i},G_{j}\in\mathcal{G}_{B}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s(\hat{\mathbf{h}}_{i},\hat{\mathbf{h}}_{j})}},$
where $\{G_{j}|y_{j}=y_{i}\}$ denotes the graphs in $\mathcal{G}_{B}$ that has the same label of $G_{i}$, and $s({\mathbf{h}}_{i},{\mathbf{h}_{j}})=\exp(cosine({\mathbf{h}}_{i},{\mathbf{h}}_{j})/\tau)$ is the similarity score between ${\mathbf{h}}_{i}$ and ${\mathbf{h}}_{j}$ based the temperature parameter $\tau$.
Also, we define a task-aware consistency loss. Note that two positive pairs $\hat{G}_{i}$ and $\tilde{G}_{i}$ are created by dropping unimportant edges from $G_{i}$. For graph classification tasks, the graph label of $\hat{G}_{i}$ and $\tilde{G}_{i}$ should remain similar to the original graph $G_{i}$. For graph regression tasks, such as counting tasks, the cycle (resp. path ) number in $\hat{G}_{i}$ and $\tilde{G}_{i}$ should be less than that in $G_{i}$, since dropping edges decreases the cycle and path numbers in $G_{i}$. Based on this analysis, we define the task-aware consistency loss $\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$ for each graph $G_{i}$ as follows:
(23) $\mathcal{L}_{cy}(\cdot)=\begin{cases}-\sum_{y\in{y}_{i}}(\log{\hat{y}^{*}}_{i}[y]+\log{\tilde{y}^{*}}_{i}[y]),&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \mathsf{ReLU}(\hat{y}^{*}_{i}-y_{i})+\mathsf{ReLU}(\tilde{y}^{*}_{i}-y_{i}),&y_{i}\in\mathbb{R}\end{cases}$
In summary, we pretrain the GNNs by the combination of contrastive learning loss and task-aware consistency loss.
$\theta^{\prime}=\arg\min_{\theta}$ $\frac{1}{B}\sum_{G_{i}\in\mathcal{G}_{B}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$
(24) $+\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$
where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss in Equation (16), and $\mathcal{L}_{cy}(\cdot)$ is the task-aware consistency loss in Equation (23). Then, we can fine-tune the GNNs with labeled data $\mathcal{LG}$ based on the task-specific loss in Equation (5).

### v2

@@ 3.2. Hierarchical Pooling-based GNNs
As mentioned in Equations (1) and (2), at the $l$-th layer, node-based GNNs learn node representations using the adjacency matrix $\mathbf{A}_{i}$ and the node hidden representations $\mathbf{H}_{i}^{(l-1)}(V_{i})$. In contrast, at the $l$-th layer, HP-based GNNs first generate a new cluster assignment matrix $\mathbf{S}_{i}^{(l)}\in\{0,1\}^{n_{l-1}\times n_{l}}$, where $n_{l}<n_{l-1}$. This matrix maps the nodes $V_{i}^{(l-1)}$ from the (l1)(l-1)-th layer to a cluster index in $\{1,\ldots,n_{l}\}$. The new adjacency matrix $\mathbf{A}_{i}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ is then reconstructed based on the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ as follows:
  • Similarity-based. Similarity-based approaches (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b) cluster nodes based on the similarity of nodes features or structures. These methods rely on predefined similarity metrics such as cosine similarity of node features, or graph partitioning algorithms to generate the cluster assignment matrix. Graclus (Mesquita et al., 2020; Dhillon et al., 2007) and CC-GNN (Li et al., 2022b) assign nodes to different clusters based on feature similarity and graph structure.
@@
3.3. Subgraph-based GNNs
@@
3.4. Graph Learning-based GNNs
Due to uncertainty and complexity in data collection, real-world graphs often contain redundant, biased, or noisy edges and features. When operating on such imperfect structures, vanilla GNNs may learn spurious correlations and thus fail to produce reliable graph representations, ultimately leading to incorrect predictions. To mitigate this issue, as shown in Figure 1 (d), recent work (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose to learn from purified or reconstructed graph structure and enhanced node features that better reflect the underlying signal to improve the quality of the learned graph representations. Given a labeled graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, the graph learning-based approaches can be formulated as bi-level optimization problem.
@@ 3.4. Graph Learning-based GNNs
$s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}\bigl(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}\bigr),$
(10) $\hfill\forall G_{i}\in\mathcal{G}$
@@
  • Preprocessing-based. Preprocessing-based approa-
    ches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly. The basic idea is to first reveal the common graph patterns and modify the graph structure and node features to recover these patterns. For example, GNN-Jaccard (Wu et al., 2019) and GNAT (Li et al., 2022a) observe that similar nodes tend to be connected by edges. Therefore, they remove edges between dissimilar nodes and add edges between similar ones. Also, GNN-SVD (Entezari et al., 2020) observes that noisy edges and node features tend to increase the rank of the adjacency matrix. Therefore, GNN-SVD reconstructs graphs by decreasing the rank of the adjacency matrix.
  • Jointly Training-based. In contrast to static preprocessing-based methods, jointly training-based approaches aim to optimize the GNN while iteratively reconstructing the graph structure and node features alongside the GNN model parameters. This iterative bi-level optimization enables the GNN to dynamically adapt the graph structure and features based on task-specific objectives. For instance, several approaches (Jin et al., 2021, 2020b; Li et al., 2024b; Franceschi et al., 2019; Sun et al., 2022; Luo et al., 2021), such as ADGNN (Li et al., 2024b), ProGNN (Jin et al., 2020b), and SimPGCN (Jin et al., 2021), propose reconstructing the edges between each pair of nodes by minimizing the GNN loss on the reconstructed structure while also reducing the rank of the adjacency matrix. Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023a), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
3.5. Self Supervised Learning-based GNNs
@@
Depending on the specific technique used to auto-generate supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main paradigms.
  • Pretext Task-based. Pretext task-based approaches  (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020b; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels. Current approaches design diverse node-level and graph-level tasks to capture intrinsic graph patterns (Wang et al., 2022b), such as predicting node attribute, node degree, and the node number. For example, HMGNN (Zang et al., 2023) pretrains GNNs by predicting the links between nodes and the number of nodes in each graph. MGSSL (Zhang et al., 2021) first masks the edges among motifs and then predicts the missing edges. MoAMa (Inae et al., 2023) first masks node features and then uses GNNs to reconstruct the node features based on the masked graphs. GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020b) predict both node attributes and edges.
  • Graph Contrastive Learning-based. Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples. The basic idea is to maximize the similarity between different views or augmentations of the same node or graph (positive pairs) while minimizing the similarity between different nodes or graphs (negative pairs). In such a way, GCL-based approaches can distinguish the similarity between nodes and graphs. In general, the SSL loss $\mathcal{L}_{ssl}(\cdot)$ in Equation (11) can be formulated as follows.
@@
    (14) $s.t.\$$\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
    where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss. Additionally, $\mathcal{L}_{positive}(\cdot)$ denotes the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views. Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs whose labels differ from $G_{i}$. One typical of the contrastive loss $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (13) can be defined based on InfoNCE loss (Zhu and others, 2021; Yeh et al., 2022) as follows:
@@
    The techniques in Equation (14) are employed to generate positive views for each node. Specifically, similarity-based methods (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes. Also, diffusion-based approaches (Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) modify the graph’s structure (e.g., by deleting or adding edges) based on global topological information, such as personalized PageRank (PPR) (Haveliwala, 1999) and motif-preserving (Sun et al., 2024). Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu and others, 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.